Circular No. 8187

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003hn IN NGC 1448 M. Salvo, M. Bessell, and B. Schmidt, Australian National University

(ANU), report that a spectrum of SN 2003hn (cf. IAUC 8186) was taken at the ANU 2.3-m telescope (+ Double Beam Spectrograph) at Siding Spring Observatory on Aug. 26.7 UT. Preliminary analysis shows very strong and broad Hα, a strong continuum in the blue, and prominent P-Cyg Hβ and Hγ lines. This is typical of a type-II supernova ≈ 2 weeks after explosion.

SUPERNOVA 2003hg IN NGC 7771 N. Elias-Rosa, S. Benetti, C. Marmo, A. Pastorello, G. Altavilla, H.

Navasardyan, M. Riello, M. Turatto and L. Zampieri, Osservatorio Astronomico di Padova; and E. Cappellaro, Osservatorio Astronomico di Capodimonte, report, from an inspection of a spectrum (range 355–780 nm; resolution 2.4 nm) obtained on Aug. 22.48 UT with the Asiago 1.8-m telescope (+ AFOSC), that SN 2003hg (IAUC 8184) is a type-II supernova caught shortly after explosion. The spectrum is dominated by Hβ and Hγ with P-Cyg profiles, while Hα still has a pure emission profile. The expansion velocity deduced from the Hβ and Hγ absorptions and from the Hα FWHM is ∼ 11300 km/s. A broad He I 587.6-nm emission is also present. A detection of a deep (EW ∼ 0.2 nm) interstellar Na I D feature suggests high reddening for SN 2003hg.

SUPERNOVA 2003hh IN UGC 12890 L. Wang, Lawrence Berkeley National Laboratory; D. Baade, Euro-

pean Southern Observatory (ESO), report on an observation of SN 2003hh (cf. IAUC 8184) with the ESO Very Large Telescopes (+ FORS1) in spectropolarimetry mode on Aug. 20. The data show that SN 2003hh is a subluminous type-Ia supernova well before optical maximum. The Si II 595.7- and 413.0-nm lines are very strong. In addition, a strong, narrow Na I D line is observed, indicating that the supernova suffers strong dust extinction from the host galaxy. The wavelength of the peak of the polarization is at 550 nm, suggesting that the dust in the host galaxy is very similar to that of the Milky Way Galaxy.

SUPERNOVA 2003ho IN ESO 235-G58 Corrigendum. On IAUC 8186, line 5, for 33′′ east read 22′′ east

2003 August 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green